OpenStudy (anonymous):

Find the equation of the line through (-6,2) and with slope -1/3 by the two following ways a) give it in point-slope form b) give it in slope-intercept form

OpenStudy (anonymous):

Point slope form: y-y_1=m(x-x_1) y-2=-1/3 (x+6). Slope Intercept form: y=mx+b, y=-1/3x + b plug in (-6,2) and solve for b, then rewrite the equation with b.